# SAGMAN: Stability Analysis of Graph Neural Networks on the Manifolds

### v3

@@ 1 Introduction
In the literature, a few studies are attempting to analyze GNN stability. Specifically, Keriven et al. first studied the stability of graph convolutional networks (GCN) on random graphs under small deformation. Later, Gama et al. and Kenlay et al. explored the robustness of various graph filters, which are then used to measure the stabilities of the corresponding (spectral-based) GNNs. However, these prior methods are limited to either synthetic graphs or specific GNN models.
In this work, we present SAGMAN, a novel framework devised to quantify the stability of GNNs through individual nodes. This is accomplished by assessing the resistance-distance distortions incurred by the map (GNN model) between low-dimensional input and output graph-based manifolds. To this end, we introduce a spectral approach for graph dimensionality reduction (GDR) to construct input/output graph-based manifolds that can well preserve effective-resistance distances between nodes. This approach allows transforming input graph data, including graph topology and node features that may originally lie in a high-dimensional space, into a low-dimensional input graph-based manifold. SAGMAN has nearly-linear algorithmic complexity and its data-centric nature allows SAGMAN to operate across various GNN variants, independent of label information, network architecture, and learned parameters, demonstrating its wide applicability. It is crucial to note that this study aims to offer significant insights for understanding and enhancing the stability of GNNs. The key technical contributions of this work are outlined below:
$\bullet$ To our knowledge, we are the first to propose a spectral method for evaluating node-level stability in GNNs. Our approach is distinguished by introducing input/output graph-based manifolds that can well preserve effective-resistance distances between nodes.
$\bullet$ We introduce a nonlinear GDR method and exploit PGMs to transform the original input graph into a low-dimensional graph-based manifold that can effectively preserve the original graph’s structural and spectral properties, enabling effective estimation of GNN stability.
$\bullet$ SAGMAN has been empirically evaluated and shown to be effective in assessing the stability of individual nodes across various GNN models in realistic graph datasets. Additionally, SAGMAN enables more potent adversarial target attacks and significantly enhances the stability of the GNNs.
$\bullet$ SAGMAN has a near-linear time complexity and is compatible with a wide range of GNN models and node label information. This makes SAGMAN highly adaptable for processing graphs of different sizes, various types of GNNs, and node-level tasks.
@@ 2 Background
2.1 Spectral Graph Theory
${A}(i,j)=\begin{cases}w(i,j)&\text{ if }(i,j)\in E\\ 0&\text{otherwise }\end{cases}\vspace{-5pt}$ (1)
@@ 2.1 Spectral Graph Theory
Lemma 2.1 is the Courant-Fischer Minimax Theorem (Golub & Van Loan, 2013) for solving the eigenvalue problem: $Lu_{k}=\lambda_{k}u_{k}$. The generalized Courant-Fischer Minimax Theorem for solving generalized eigenvalue problem $L_{X}v_{k}=\lambda_{k}L_{Y}v_{k}$ can be expressed as follows:
Lemma 2.2.
(The Generalized Courant-Fischer Minimax Theorem) Given two Laplacian matrices $L_{X},L_{Y}\in\mathbb{R}^{|V|\times|V|}$ such that $null(L_{Y})\subseteq null(L_{X})$, the $k$-th largest eigenvalue of $L_{Y}^{+}L_{X}$ can be computed under the condition of $1\leq k\leq rank(L_{Y})$ by:
2.2 Stability Analysis of ML Models on the Manifolds
The stability of an ML model refers to the ability of the model to produce consistent output despite small variations or noise in the input (Szegedy et al., 2013). Let $M$ denote an ML model, which operates on input $X$ to yield output $Y$, i.e., $Y=M(X)$. The stability of the ML model can be assessed by evaluating the distance distortions incurred by the maps between low-dimensional input and output manifolds. Specifically, a distance mapping distortion (DMD) metric has been introduced to evaluate the distance distortions between graph-based manifolds (Cheng et al., 2021). For two input data samples $p$ and $q$, the DMD metric denoted by $\delta^{M}(p,q)$ is defined as the ratio of the distance $d_{Y}(p,q)$ measured on the output graph-based manifold $G_{Y}=(V,E_{Y})$ to the one $d_{X}(p,q)$ on the input manifold $G_{X}=(V,E_{X})$ (Cheng et al., 2021):
$DMD=\delta^{M}(p,q)\overset{\mathrm{def}}{=}\frac{d_{Y}(p,q)}{d_{X}(p,q)}.$ (4)
By evaluating DMD for each pair of data samples, it becomes possible to assess the stability of the ML model: when two nearby data samples on the input manifold are mapped to two distant ones on the output manifold, it will imply a large local Lipschitz constant and thus poor stability of the ML model near these data samples.
2.3 Stability of GNNs
The stability of a GNN refers to its performance (output) stability in the presence of edge/node perturbations (Sun et al., 2020). This includes the ability to maintain the fidelity of predictions and outcomes when subjected to changes such as edge alterations or feature attacks. To elucidate, consider the scenario where the input graph data is subject to minor modifications, such as the addition or removal of a node, a change in node features, or a slight alteration in the graph structure. A desired GNN model is expected to exhibit good stability, wherein every predicted output or the graph embeddings do not change drastically in response to the aforementioned minor perturbations (Jin et al., 2020; Zhu et al., 2019). On the contrary, an unstable GNN would exhibit a considerable change in part of outputs or embeddings, potentially leading to erroneous inferences or predictions.
The importance of analyzing the stability of GNNs has been underlined by several recent studies. For instance, the vulnerability of GNNs to adversarial attacks has been studied in which small perturbations on targeted nodes can result in significant misclassifications (Zügner et al., 2018). In addition, although recent studies have investigated the stability issues of GNNs on synthetic graphs or specific models, they failed to develop a unified framework for evaluating the stability of GNNs.  (Keriven et al., 2020; Gama et al., 2020; Kenlay et al., 2021).
3 SAGMAN for Stability Analysis of GNNs
3.1 GNN Stability Analysis with DMDs is Nontrivial
When adopting the DMD metric for the stability analysis of GNN models, a natural question that arises is whether we can directly use the input graph structure of a GNN model as the input graph-based manifold for DMD calculations. Unfortunately, as shown in our empirical results reported in Table 2 and Appendix C, directly using the input graph topology for DMD evaluations does not lead to satisfactory stability analysis results. This is due to the fact that the original input graph data may not always lie near a low-dimensional space (Bruna et al., 2013), while meaningful DMD-based stability analysis requires that both input and output data samples lie near low-dimensional manifolds.
3.2 Overview of SAGMAN
Figure 1: The proposed three-phase SAGMAN framework for stability analysis of GNNs on the manifolds.
To extend the applicability of the DMD metric to the GNN settings, we introduce a spectral framework, SAGMAN, for stability analysis of GNNs. A key ingredient of SAGMAN is a novel distance-preserving GDR algorithm for converting the original input graph data (including node features and graph topology) that may lie in high-dimensional space into a low-dimensional graph-based manifold. As shown in Figure 1, the proposed SAGMAN framework consists of three main phases: Phase 1 for creating the input graph embedding matrix (based on both node features and spectral graph properties) indispensable to the subsequent GDR step, Phase 2 for constructing low-dimensional input and output graph-based manifolds exploiting a probabilistic graphical model (PGM) approach, and Phase 3 for DMD-based node stability evaluation leveraging a spectral graph embedding scheme using generalized Laplacian eigenvalues/eigenvectors. A detailed exposition of our framework is presented in the following sections.
3.3 Phase 1: Embedding Matrix Construction for GDR
GDR via Laplacian Eigenmaps.
In this work, we exploit spectral graph embedding and the widely recognized nonlinear dimensionality reduction algorithm, Laplacian Eigenmaps (Belkin & Niyogi, 2003), for reducing the dimensionality of an input graph. The algorithm flow of Laplacian Eigenmaps starts with constructing an undirected (nearest-neighbor) graph where each node represents a (high-dimensional) data sample (vector) and each edge encodes the similarity between two data samples. Then the eigenvectors corresponding to the smallest eigenvalues of the graph Laplacian matrix are exploited for mapping each node (sample) into a low-dimensional space for preserving the local relationships between data samples.
Spectral Embedding with Eigengaps.
To apply Laplacian Eigenmaps for dimensionality reduction of a given graph with $N$ nodes, a straightforward approach is to first compute a $N\times N$ spectral embedding matrix using the complete set of graph Laplacian eigenvectors/eigenvalues (Ng et al., 2001), such that each node can be represented by an $N$-dimensional vector; next Laplacian Eigenmap will take this embedding matrix as input for the subsequent steps. However, computing the full set of eigenvalues/eigenvectors will be too expensive for large graphs. To address this complexity issue, a recent theoretical result on spectral graph clustering  (Peng et al., 2015) will be exploited in SAGMAN, which shows that the existence of a significant gap between two consecutive eigenvalues $\Upsilon(k)=\frac{\lambda_{k+1}}{\rho(k)}$, where $\rho(k)$ denotes the $k$-way expansion constant and $\lambda_{k+1}$ denotes the $(k+1)$-th smallest eigenvalue of the normalized Laplacian matrix, implies the existence of a $k$-way partition for which every cluster has low conductance, and that the underlying graph is a well-clustered graph.
Definition 3.1.
For a connected graph $G=(V,E,w)$ with its $k$ smallest nonzero Laplacian eigenvalues denoted by $0<\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{k}$ and their corresponding eigenvectors denoted by $u_{1},u_{2},\dots,u_{k}$, its weighted spectral embedding matrix is defined as $U_{k}=\left[\frac{u_{1}}{\sqrt{\lambda_{1}}},\dots,\frac{u_{k}}{\sqrt{\lambda_{k}}}\right]\in{\mathbb{R}^{|V|\times k}}$.
For a graph with a significant eigengap $\Upsilon(k)$, the above embedding matrix allows representing each node with a $k$-dimensional vector such that the effective-resistance distance between arbitrary node pairs can be well approximated by $d^{eff}({p,q})\approx|U_{k}^{\top}e_{p,q}|^{2}_{2}$, where ${e_{p}}\in\mathbb{R}^{|V|}$ denotes the standard basis vector with all zero entries except for the $p$-th entry being $1$, and ${e_{p,q}}=e_{p}-e_{q}$.
Eigengap for Graph Dimension Estimation.
Graph dimension is defined as the minimum integer that permits a ‘classical representation’ of the graph within an Euclidean space ($\mathbb{R}^{n}$) while preserving the unit length of all edges (Erdös et al., 1965). However, the precise quantification of either the graph dimension or the Euclidean dimension is computationally challenging, being an NP-hard problem (Schaefer, 2012). Fortunately, the existence of a significant eigengap $\Upsilon(k)$ implies the original graph can be very well represented in a $k$-dimensional space (Peng et al., 2015), and thus the graph dimensionality is approximately $k$. While precisely quantifying $\Upsilon(k)$ still remains challenging in practice, SAGMAN circumvents the need for the exact computation of graph dimensions. Instead, the identified eigengap acts as an indicative marker for assessing the suitability of SAGMAN for a specific graph: the ones with significant eigengaps will be more suitable for the SAGMAN framework since they can be well represented in low-dimensional space. Empirically, for datasets with the class number $c$, the value of $k$ can be approximated as $k\approx 10c$ to capture significant eigengaps effectively (Deng et al., 2022).
3.4 Phase 2: Manifold Construction via PGMs
While the original algorithm in Laplacian Eigenmaps suggests creating graph-based manifolds using $\epsilon$- or $k$-nearest-neighbor graphs, it is unclear whether the resulting graph-based manifolds can well preserve the effective-resistance distances on the original graph.
PGM for Graph Topology Learning from Data.
PGMs, also known as Markov Random Fields (MRFs), are powerful tools in machine learning and statistical physics for representing complex systems with intricate dependency structures (Roy et al., 2009). PGMs encode the conditional dependencies between random variables through an undirected graph structure. Recent results show that the graph structure learned through PGM will have its resistance distances encoding the Euclidean distances between their corresponding data samples  (Feng, 2021). Since each column vector in the embedding matrix $U_{k}$ in Definition 3.1 corresponds to a data sample for graph topology learning, the low-dimensional manifold constructed via PGM will well preserve the resistance distances in the original graph. However, the state-of-the-art method may require numerous iterations to achieve convergence (Feng, 2021), which limits its applicability in learning large graphs.
PGM via Spectral Sparsification.
In the proposed SAGMAN framework, we exploit PGMs for creating low-dimensional input graph-based manifolds using the embedding matrix $U_{k}$ in Definition 3.1, and the output manifold using GNN’s post-softmax vectors as shown in Figure 1. In the following, we provide a detailed description for constructing the input manifold, while the output one can be created in a similar way. For a given input embedding matrix $X=U_{k}\in{\mathbb{R}^{|V|\times k}}$, the maximum likelihood estimation (MLE) of the precision matrix $\Theta$ (PGM) can be obtained by solving the following convex problem (Dong et al., 2019):
${\max_{\Theta}}:F(\Theta)=\log\det(\Theta)-\frac{1}{k}Tr({X^{\top}\Theta X})$ (5)
where $\Theta={L}+\frac{1}{\sigma^{2}}I$, $Tr(\bullet)$ denotes the matrix trace, ${L}$ denotes the set of valid Laplacian matrices, $I$ denotes the identity matrix, and $\sigma^{2}>0$ denotes prior feature variance. Expanding the Laplacian matrix of $G=(V,E,w)$ with $L=\sum\limits_{\left({p,q}\right)\in E}{w_{p,q}}e_{p,q}e^{\top}_{p,q}$, where $w_{p,q}=\frac{1}{\|X^{\top}e_{p,q}\|_{2}^{2}}$ denotes the weight of the edge $(p,q)$, allows for expressing the objective function in Equation 5 as $F=F_{1}-\frac{1}{k}F_{2}$, where:
$\begin{split}F_{1}&=\log\det(\Theta)=\sum\limits_{i=1}^{|V|}\log(\lambda_{i}+\frac{1}{\sigma^{2}})\\ F_{2}&=Tr({X^{\top}\Theta X})=\frac{Tr(X^{\top}X)}{\sigma^{2}}+\sum\limits_{\left({p,q}\right)\in E}{w_{p,q}}\|X^{\top}e_{p,q}\|_{2}^{2}.\end{split}$ (6)
Taking partial derivatives with respect to $w_{p,q}$ leads to (Zhang et al., 2022):
$\begin{split}\frac{\partial F_{1}}{\partial w_{p,q}}&=\sum\limits_{i=1}^{|V|}\frac{1}{\lambda_{i}+1/\sigma^{2}}\frac{\partial\lambda_{i}}{\partial w_{p,q}}=d^{eff}({p,q})\\ \frac{\partial F_{2}}{\partial w_{p,q}}&=\|X^{\top}e_{p,q}\|_{2}^{2}=d^{dat}({p,q})=\frac{1}{w_{p,q}},\end{split}$ (7)
where $d^{eff}({p,q})$ and $d^{dat}({p,q})$ denote the effective-resistance distance and $\ell_{2}$ distance between nodes $p$ and $q$, respectively. To maximize the objective function $F$, an effective approach is to aggressively prune an initial graph (e.g. a nearest-neighbor graph) according to the gradients in Equation 7 by removing the edges with small resistance distance $d^{eff}({p,q})$ but large $\ell_{2}$ distance $d^{dat}({p,q})$ between nodes $p$ and $q$. This pruning strategy is equivalent to pruning edges with small distance ratios, defined as:
$\rho_{p,q}=\frac{d^{eff}({p,q})}{d^{dat}({p,q})}=w_{p,q}d^{eff}({p,q}).$ (8)
Since $\rho_{p,q}$ is exactly the edge sampling probability for spectral graph sparsification (Spielman & Teng, 2011), the proposed edge pruning strategy is equivalent to pruning the initial graph through spectral sparsification.
Spectral Sparsification via Graph Decomposition
To compute the edge sampling probability $\rho_{p,q}$ for each edge ($p,q$) requires solving Laplacian matrices multiple times (Spielman & Srivastava, 2008), which makes the original sparsification method computationally costly for large problems. An alternative approach for spectral sparsification employs a short-cycle graph decomposition scheme (Lemma 3.2) (Chu et al., 2020), which partitions an unweighted graph $G$ into multiple disjoint cycles by removing a fixed number of edges while ensuring a bound on the length of each cycle. In the last, the algorithm combines short-cycle decomposition with low-stretch spanning trees (LSSTs) to construct the sparsified graph for preserving the spectral properties of the original graph (Liu et al., 2019). However, such methods can only handle unweighted graphs.
Lemma 3.2.
@@ 3.4 Phase 2: Manifold Construction via PGMs
Spectral sparsification of an undirected graph $G$ with its Laplacian denoted by $L_{G}$ can be obtained by leveraging a short-cycle decomposition algorithm, which returns a sparsified graph $H$ with its Laplacian denoted by $L_{H}$ such that for all real vectors $x$, $x^{\top}L_{G}x\approx x^{\top}L_{H}x$ (Chu et al., 2020).
As a substantial extension of prior short-cycle-based algorithms, we introduce an improved spectral sparsification algorithm, as shown in Figure 2. Our approach utilizes a low-resistance-diameter (LRD) decomposition scheme to limit the length of each cycle as measured by the effective-resistance metric. This method is particularly effective for sparsifying weighted graphs. The key idea in our method is to efficiently compute the effective resistance of each edge (see details in Appendix D) and leverage a multi-level framework to decompose the graph into several disjoint cycles bounded by an effective-resistance threshold. As shown in Figure 2, the subgraph corresponding to each node cluster will be sparsified into an LSST, while inter-cluster (bridge-like) edges will be retained to preserve the original structural properties. It is worth noting that these inter-cluster edges can be inserted into the original graph to significantly enhance the stability of GNN models, as shown in Table 4.

### v4

@@ 1 Introduction
In the literature, while there are studies primarily focused on developing more stable GNN architectures (Wu et al., 2023; Zhao et al., 2024; Song et al., 2022; Gravina et al., 2022), a few attempts to analyze GNN stability. Specifically, (Keriven et al., 2020) first studied the stability of graph convolutional networks (GCN) on random graphs under small deformation. Later, (Gama et al., 2020) and (Kenlay et al., 2021) explored the robustness of various graph filters, which are then used to measure the stabilities of the corresponding (spectral-based) GNNs. However, these prior methods are limited to either synthetic graphs or specific GNN models.
In this work, we present SAGMAN, a novel framework devised to quantify the stability of GNNs through individual nodes. This is accomplished by assessing the resistance-distance distortions incurred by the nonlinear map (GNN model) between low-dimensional input and output graph-based manifolds. It is crucial to note that this study aims to offer significant insights for understanding and enhancing the stability of GNNs. The key technical contributions of this work are outlined below:
$\bullet$ This study introduces a spectral framework (SAGMAN) for measuring the stability of GNN models at the node level. This is achieved by effectively assessing the distance distortions caused by the maps between the input and output smooth manifolds.
$\bullet$ To construct the input smooth manifold for stability analysis in SAGMAN, we introduce a nonlinear graph dimension reduction (GDR) framework to transform the original input graph (with node features) into a low-dimensional graph-based manifold that can well preserve the original graph’s spectral properties, such as effective resistance distances between nodes.
$\bullet$ SAGMAN has been empirically evaluated and shown to be effective in assessing the stability of individual nodes across various GNN models in realistic graph datasets. Moreover, SAGMAN allows for more powerful adversarial targeting attacks and greatly improving the stability (robustness) of the GNNs.
$\bullet$ SAGMAN has a near-linear time complexity and its data-centric nature allows it to operate across various GNN variants, independent of label information, network architecture, and learned parameters, demonstrating its wide applicability.
@@ 2 Background
2.1 Stability Analysis of ML Models on the Manifolds
The stability of a machine learning (ML) model refers to its ability to produce consistent outputs despite small variations or noise in the input data (Szegedy et al., 2013). To assess this stability, we utilize the Distance Mapping Distortion (DMD) metric (Cheng et al., 2021). For two input data samples $p$ and $q$, the DMD metric $\delta^{M}(p,q)$ is defined as the ratio of their distance on the output manifold to the one on the input manifold:
$\delta^{M}(p,q)\overset{\mathrm{def}}{=}\frac{d_{Y}(p,q)}{d_{X}(p,q)}.$ (1)
By evaluating $\delta^{M}(p,q)$ for each pair of data samples, we can assess the stability of the ML model. Specifically, if two nearby data samples on the input manifold are mapped to distant points on the output manifold, this indicates a large $\delta^{M}(p,q)$ or equivalently a large local Lipschitz constant, and thus poor stability of the model near these samples; On the other hand, a small $\delta^{M}(p,q)$ implies that the model is stable in that region.
2.2 Previous Investigations on the Stability of GNNs
The stability of a GNN refers to its output stability in the presence of edge/node perturbations (Sun et al., 2020). This includes maintaining the fidelity of predictions and outcomes when subjected to changes such as edge alterations or feature attacks. A desired GNN model is expected to exhibit good stability, wherein every predicted output or the graph embeddings do not change drastically in response to the aforementioned minor perturbations (Jin et al., 2020; Zhu et al., 2019). Several recent studies have underlined the importance of analyzing the stability of GNNs. For instance, Sharma et al. (2023) examined the task and model-agnostic vulnerabilities of GNNs, demonstrating that these networks remain susceptible to adversarial perturbations regardless of the specific downstream tasks. Furthermore, Huang et al. (2023) investigated robust graph representation learning via predictive coding, proposing a method that enhances GNN stability by reconstructing input data to mitigate the effects of adversarial attacks.
Moreover, while recent studies have investigated the stability issues of Graph Neural Networks (GNNs) on synthetic graphs or specific models (Keriven et al., 2020; Kenlay et al., 2021), they have not provided a unified framework for evaluating GNN stability. The latest state-of-the-art method (Li et al., 2024) develops such a unified framework but does not consider feature perturbations. This omission leaves a significant gap in understanding how GNNs respond to changes in node features, which is crucial for applications where feature data may be noisy or subject to perturbations.
3 The SAGMAN Framework for Stability Analysis of GNNs
Challenges in Applying DMD Metrics to GNN Stability Analysis.
When adopting the DMD metric for the stability analysis of GNN models, a natural approach is to use graph-based manifolds (details in Appendix A.1) for calculating DMDs. However, directly using the input graph structure as the input graph-based manifold may not produce satisfactory results as shown in our empirical results presented in Table 2 and Appendix A.8. This inadequacy stems from the fact that the original input graph data (including node features) may not reside near a low-dimensional manifold (Bruna et al., 2013), while meaningful DMD-based stability analysis requires both the input and output data samples to lie near low-dimensional manifolds (Cheng et al., 2021). Therefore, a naive application of DMD metrics on the original graph structure is insufficient for assessing the stability of GNNs.
3.1 Overview of SAGMAN
Figure 1: The proposed SAGMAN framework for stability analysis of GNNs on the manifolds.
To extend the applicability of the DMD metric to GNN settings, we introduce SAGMAN, a spectral framework for stability analysis of GNNs. A key component of SAGMAN is a novel distance-preserving Graph Dimensionality Reduction (GDR) algorithm that transforms the original input graph data—including node features and graph topology, which may reside in high-dimensional space—into a low-dimensional graph-based manifold.
As illustrated in Figure 1, the SAGMAN framework comprises three main phases:
  • Phase 1: Creation of the input graph embedding matrix based on both node features and spectral properties of the graph. This embedding matrix is essential for the subsequent GDR step.
  • Phase 2: Construction of low-dimensional input and output graph-based manifolds using a Probabilistic Graphical Model (PGM) approach.
  • Phase 3: Node stability evaluation using the DMD metric, leveraging a spectral graph embedding scheme that utilizes generalized Laplacian eigenvalues and eigenvectors.
Detailed descriptions of each phase are provided in the following sections. The complete algorithmic flow of SAGMAN is shown in Section 3.5.
3.2 Phase 1: Embedding Matrix Construction for GDR
The calculation of the DMD metric inherently relies on pairwise distances between nodes. For graph-based manifolds, two widely used metrics for evaluating pairwise node distances are: (1) the shortest-path distance and (2) the effective resistance distance. As discussed in Appendix A.1, the effective resistance distance is more closely related to the graph’s structural (spectral) properties, providing a more meaningful measure of connectivity between nodes. Moreover, previous studies have demonstrated a significant correlation between effective resistance distances and the stability of machine learning models (Cheng et al., 2021). Therefore, our focus is on performing dimensionality reduction on the input graph while preserving its original effective resistance distances.
Graph Dimensionality Reduction via Laplacian Eigenmaps.
To achieve dimensionality reduction of the input graph, we utilize the widely recognized nonlinear dimensionality reduction algorithm, Laplacian Eigenmaps (Belkin & Niyogi, 2003). The Laplacian Eigenmaps algorithm begins by constructing an undirected graph where each node represents a high-dimensional data sample, and edges encode the similarities between data samples. It then computes the eigenvectors corresponding to the smallest eigenvalues of the graph Laplacian matrix, which are used to map each node into a low-dimensional space while preserving the local relationships between the data samples.
Spectral Embedding with Eigengaps.
For a graph with $N$ nodes, a straightforward approach to apply Laplacian Eigenmaps is to compute an $N\times N$ spectral embedding matrix using the complete set of graph Laplacian eigenvectors and eigenvalues (Ng et al., 2001), representing each node with an $N$-dimensional vector. However, computing the full set of eigenvalues and eigenvectors is computationally prohibitive for large graphs.
To address this issue, we leverage a theoretical result from spectral graph clustering (Peng et al., 2015), which shows that the existence of a significant gap between consecutive eigenvalues—known as an eigengap—implies that the graph can be well represented in a lower-dimensional space. Specifically, the eigengap is defined as $\Upsilon(k)=\frac{\lambda_{k+1}}{\rho(k)}$, where $\rho(k)$ denotes the $k$-way expansion constant, and $\lambda_{k+1}$ is the $(k+1)$-th smallest eigenvalue of the normalized Laplacian matrix. A significant eigengap indicates the existence of a $k$-way partition where each cluster has low conductance, meaning the graph is well-clustered. Based on this, we can approximate the spectral embedding using only the first $k$ eigenvalues and eigenvectors. We define the weighted spectral embedding matrix as follows:
Definition 3.1.
For a connected graph $G=(V,E,w)$ with its $k$ smallest nonzero Laplacian eigenvalues denoted by $0<\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{k}$ and corresponding eigenvectors $u_{1},u_{2},\dots,u_{k}$, the weighted spectral embedding matrix is defined as $U_{k}=\left[\frac{u_{1}}{\sqrt{\lambda_{1}}},\dots,\frac{u_{k}}{\sqrt{\lambda_{k}}}\right]\in\mathbb{R}^{|V|\times k}$.
For a graph with a significant eigengap $\Upsilon(k)$, this embedding matrix allows us to represent each node with a $k$-dimensional vector such that the effective resistance distance between any pair of nodes can be well approximated by $d^{\mathrm{eff}}(p,q)\approx\|U_{k}^{\top}e_{p,q}\|_{2}^{2}$, where $e_{p}\in\mathbb{R}^{|V|}$ is the standard basis vector with a 1 at position $p$ and zeros elsewhere, and $e_{p,q}=e_{p}-e_{q}$.
Using Eigengaps for Graph Dimension Estimation.
Determining the precise graph dimension required to embed a graph into Euclidean space while preserving certain properties (e.g., unit edge lengths) is an NP-hard problem (Erdös et al., 1965; Schaefer, 2012). However, the presence of a significant eigengap $\Upsilon(k)$ suggests that the graph can be well represented in a $k$-dimensional space (Peng et al., 2015), making $k$ an approximate measure of the graph’s intrinsic dimensionality. While computing the exact value of $\Upsilon(k)$ may be challenging in practice, we can use the identified eigengap as an indicator of the suitability of SAGMAN for a given graph: graphs with significant eigengaps are more suitable to our framework since they can be effectively represented in low-dimensional spaces. Empirically, for datasets with $c$ classes, we can approximate $k$ as $k\approx 10c$ to effectively capture significant eigengaps (Deng et al., 2022).
3.3 Phase 2: Manifold Construction via PGMs
While the original Laplacian Eigenmaps algorithm suggests constructing graph-based manifolds using $k$-nearest-neighbor graphs, we find that these manifolds do not adequately preserve the original effective resistance distances. Consequently, they are not suitable for GNN stability analysis, as empirically demonstrated in Appendix A.6.
Probabilistic Graphical Models (PGMs) for Graph-based Manifold Learning.
PGMs, also known as Markov Random Fields (MRFs), are powerful tools in machine learning and statistical physics for representing complex systems with intricate dependency structures (Roy et al., 2009). PGMs encode the conditional dependencies between random variables through an undirected graph structure (see Appendix A.3 for more details). Recent studies have shown that the graph structure learned through PGMs can have resistance distances that encode the Euclidean distances between their corresponding data samples (Feng, 2021).
In our context, each column vector in the embedding matrix $U_{k}$ (as defined in Definition 3.1) corresponds to a data sample used for graph topology learning. By constructing the low-dimensional graph-based manifold via PGMs, we can effectively preserve the resistance distances from the original graph. Empirical evidence, detailed in Appendix A.6, supports that preserving these distances is essential for distinguishing between stable and unstable nodes. Therefore, our methodology leverages PGMs to maintain accurate effective resistance distances.
However, existing methods for learning PGMs may require numerous iterations to achieve convergence (Feng, 2021), limiting their applicability to large graphs.
Scalable PGM via Spectral Sparsification.
In the proposed SAGMAN framework, we employ PGMs to create low-dimensional input graph-based manifold $G_{X}=(V,E_{X})$ using the embedding matrix $U_{k}$ from Definition 3.1, and output manifold $G_{Y}=(V,E_{Y})$ using the GNN’s post-softmax vectors, as illustrated in Figure 1. Below, we detail the construction of the input manifold; the output manifold can be constructed similarly. Given the input embedding matrix $X=U_{k}\in\mathbb{R}^{|V|\times k}$, the maximum likelihood estimation (MLE) of the precision matrix $\Theta$ (PGM) can be obtained by solving the following convex optimization problem (see Appendix A.3 for more details) (Dong et al., 2019):
$\max_{\Theta}\quad F(\Theta)=\log\det(\Theta)-\frac{1}{k}\operatorname{Tr}(X^{\top}\Theta X),$ (2)
where $\Theta=\mathcal{L}+\frac{1}{\sigma^{2}}I$, $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix, $\mathcal{L}$ is a valid Laplacian matrix, $I$ is the identity matrix, and $\sigma^{2}>0$ is a prior feature variance. To solve this, we give the following theorem:
Theorem 3.2.
Maximizing the objective function in Equation 2 can be achieved in nearly-linear time via the following edge pruning strategy equivalent to spectral sparsification of the initial dense nearest-neighbor graph. Specifically, edges with small distance ratios
$\rho_{p,q}=\frac{d^{\mathrm{eff}}(p,q)}{d^{\mathrm{dat}}(p,q)}=w_{p,q}\,d^{\mathrm{eff}}(p,q)$
are pruned, where $d^{\mathrm{eff}}(p,q)$ is the effective resistance distance between nodes $p$ and $q$, $d^{\mathrm{dat}}(p,q)=\|X_{p}-X_{q}\|_{2}^{2}$ is the data distance between the embeddings of nodes $p$ and $q$, and $w_{p,q}=\dfrac{1}{d^{\mathrm{dat}}(p,q)}$ is the weight of edge $(p,q)$.
The proof for Theorem 3.2 is available in Appendix A.4.
Spectral Sparsification via Graph Decomposition.
Computing the edge sampling probability $\rho_{p,q}$ for each edge $(p,q)$ requires solving Laplacian matrices multiple times (Spielman & Srivastava, 2008), making the original sparsification method computationally expensive for large graphs. An alternative approach employs a short-cycle graph decomposition scheme (Chu et al., 2020), which partitions an unweighted graph $G$ into multiple disjoint cycles by removing a fixed number of edges while ensuring a bound on the length of each cycle. However, such methods are limited to unweighted graphs.
@@ 3.3 Phase 2: Manifold Construction via PGMs
Spectral sparsification of an undirected graph $G$, with Laplacian $L_{G}$, can be achieved by leveraging a short-cycle decomposition algorithm that returns a sparsified graph $H$, with Laplacian $L_{H}$, such that for all real vectors $x$, $x^{\top}L_{G}x\approx x^{\top}L_{H}x$ (Chu et al., 2020).
To extend these methods to weighted graphs, we introduce an improved spectral sparsification algorithm, illustrated in Figure 2. Our approach utilizes a low-resistance-diameter (LRD) decomposition scheme to limit the length of each cycle as measured by the effective resistance metric. This method is particularly effective for sparsifying weighted graphs.
@@
The key idea is to efficiently compute the effective resistance of each edge (see Appendix A.9 for details) and employ a multilevel framework to decompose the graph into several disjoint clusters bounded by an effective resistance threshold. Importantly, the inter-cluster edges identified during this process can be inserted back into the original graph to significantly enhance the stability of GNN models, as demonstrated in Section 4.5.